OpenStudy (anonymous):

Input: 1, 2, 3, 4, 5 Output: 1, 3, 7, 15, 31 Find the pattern equation

OpenStudy (anonymous):

x+1=2x+1

OpenStudy (anonymous):

\[y _{n} = y _{n-1}+2^{x _{n-1}}\]

OpenStudy (anonymous):

you lost me on that one. Can you explain?

OpenStudy (anonymous):

y (2) = 3 = y(0)+2^(x(0)) y(0) = 1. x(0) = 1 3 = 1+2^1

OpenStudy (anonymous):

y(3) = 7 = y(1)+2^x(1) y(1) = 3, x(1) = 2 7 = 3+2^(2) = 3+4

OpenStudy (anonymous):

what if x =120? how would you solve that using the equation?

OpenStudy (anonymous):

oh man i was really wrong

OpenStudy (anonymous):

if x=120 you would get something like 1.3x10^36